The Chocolate Baker specializes in chocolate baked goods. The firm has long assessed the profitability of a product line by comparing revenues to the cost of goods sold. However, Barry Love, the firm’s new accountant, wants to use an activity-based costing system that takes into consideration the cost of the delivery person. Listed below are activity and cost information relating to two of Chocolate Baker’s major products.

Muffins

Cheesecake

Revenue

$53,000

$46,000

Cost of goods sold

26,000

21,000

Delivery activity:

Number of deliveries

150

85

Average length of delivery

10 minutes

15 minutes

Cost per hour for delivery

$20.00

$20.00

Using activity-based costing, which one of the following statements is correct?

A. The cheesecakes are $75 more profitable.

B. The muffins have a higher profitability as a percentage of sales and therefore are more advantageous.

C. The muffins are $2,000 more profitable.

D. The muffins are $1,925 more profitable.
Answer #1
Cost of Delivery:
Muffins 500 =150*20*(10/60)
Cheesecake 425 =85*20*(15/60)
Muffins Cheesecake
Revenue 53000 46000
Less: Cost of goods sold 26000 21000
Less: Delivery cost 500 425
Profit 26500 24575
Difference 1925 =26500-24575
The muffins are $1,925 more profitable.
Option D is correct